A line passes through (–10, 5) and (–9, 10).
a. Write an equation for the line in point-slope form.
b. Rewrite the equation in standard form using integers.
Plz help me ive been stuck with this problem 4 a long time.

Answer by stanbon(75887)

A line passes through (–10, 5) and (–9, 10).
a. Write an equation for the line in point-slope form.
slope = (10-5)/(-9--10) = 5/1 = 5
---
Form: y = mx + b
You know y = 5 when x = -10 and you knoe m = 5.
So solve for "b":
5 = 5(-10)+b
b = 55
Equation:
y = 5x+55
------------------------------
b. Rewrite the equation in standard form using integers.
Standard form: ax + by = c
5x -y = -55
==================
cheers,
Stan H.